RTYLQ:

I have a question about the two-step equation process. It's been a while since I've had to do these. Can someone work me through how to solve this equation? 17 = 5k - 2

Timmyspu:

Well you have to use the y=mx+b equation and AZ correct me if I am wrong 17=5k-2 first we have to get rid of the -2 17=5k-2 -2= -2 15=5k Now we have to get rid of the 5 from k 15=5k /5 /5 3=k

AZ:

it's official

Timmyspu:

If you would like to further explain the problem better your welcome to AZ.

AZ:

Timmyspu is incorrect

RTYLQ:

@timmyspu wrote:
Well you have to use the y=mx+b equation and AZ correct me if I am wrong 17=5k-2 first we have to get rid of the -2 17=5k-2 -2= -2 15=5k Now we have to get rid of the 5 from k 15=5k /5 /5 3=k
Thank you lots! I understand it much much better now!

sammixboo:

@rtylq wrote:
@timmyspu wrote:
Well you have to use the y=mx+b equation and AZ correct me if I am wrong 17=5k-2 first we have to get rid of the -2 17=5k-2 -2= -2 15=5k Now we have to get rid of the 5 from k 15=5k /5 /5 3=k
Thank you lots! I understand it much much better now!
Timmy is not correct about this, you can check and replace k with 3 to double check. This problem has nothing to do with the slope equation

AZ:

17 = 5k - 2 We want to move all the numbers to one side and get 'k' all by itself You can see that we're subtracting 2 from 5k 5k MINUS 2 to get the 2 on the other side so that we have 5k on one side, we have to do the opposite of subtraction to that 2 what is the opposite of subtraction? is it addition? multiplication? division?

Timmyspu:

Thank you for correcting me I apologize for doing it wrong I was thinking it was a slope intercept question but I need to brush up on my math.